OpenStudy (anonymous):

0.24 radians is how many degrees?

OpenStudy (anonymous):

\[(.24/\pi )180=13.751{}^{\circ} \]

myininaya (myininaya):

\[\frac{24}{100} \cdot \frac{180^o}{\pi}=\frac{6}{25} \cdot \frac{180^o}{\pi}=(\frac{6 \cdot 180}{25 \pi})^o\]

myininaya (myininaya):

\[=(\frac{6 \cdot 36}{5\pi})^o\]

myininaya (myininaya):

\[=(\frac{216}{5\pi})^o\]

OpenStudy (anonymous):

great